# Neural Stereo Video Compression with Hybrid Disparity Compensation

### v1

@@ Appendix C Model complexity
Ours 16.39 536.90
Table S4: Model complexity comparison. * denotes our reproduced baseline model.
@@
Table S4 compares the FLOPs and parameter counts of our model with DCVC-TCM [50], DCVC-TCM*, LSVC [11], and LLSS [25], using a 512$\times$512 input frame pair. The FLOPs and parameter values for LLSS are taken from its published work, while LSVC values were provided by the authors. We measured FLOPs and parameters for DCVC-TCM, DCVC-TCM*, and our method on an NVIDIA RTX 4090 GPU.
@@
As shown in Table S4, DCVC-TCM*, a modified version of DCVC-TCM (our backbone model), reduces FLOPs from 735G to 431G, significantly lowering computational complexity. Integrating the proposed HDC-FER and HDC-EM modules into DCVC-TCM* slightly increases FLOPs to 536.90G while still maintaining a substantially lower computational cost than DCVC-TCM. Compared to previous SOTA NSVC methods LSVC and LLSS, our model achieves superior compression performance with significantly reduced computational overhead. These results highlight the computational efficiency and effectiveness of HDC-FER and HDC-EM, making our method suitable for deployment in resource-constrained environments.
Appendix D Network Structure
Our network consists of five key components: motion estimation, motion compression, motion compensation, context compression, and frame reconstruction. For motion estimation, we adopt a lightweight version of RAFT [54], initialized with official pretrained weights. The motion compensation and frame reconstruction modules follow the structures of DEVC-TCM [50], with a reduced number of channels to lower model complexity.
We further detail the motion and context compression modules, where our proposed HDC-FER and HDC-EM modules are integrated. Specifically, HDC-FER is incorporated into the encoder, decoder, hyper-encoder, and hyper-decoder of both motion and context compression networks. Meanwhile, HDC-EM acts as the entropy model, performing entropy prediction for both modules. The structures of the motion and context compression networks are illustrated in Fig. S6.
Figure S6: The structure of motion and context compression network

### v2

@@ IV-C Comparison with State-of-the-art Methods
Ours 16.39 536.90 8.33fps 10.00fps
Rate-Distortion Performance. The rate–distortion (RD) curves in Fig. 4 compare our proposed method with baseline approaches in terms of Bpp–PSNR across four datasets: KITTI 2012 [19], KITTI 2015 [20], and two general-scene video sequences, Kendo and Balloons, from the Nagoya dataset [21]. As shown in the figure, our method consistently achieves state-of-the-art performance across all datasets, demonstrating substantial improvements in RD efficiency over both traditional and neural-based baselines.
Table I presents a detailed quantitative analysis using BD-Rate [61] as the evaluation metric, computed based on Bpp–PSNR with MV-HEVC [2] serving as the anchor. On autonomous driving datasets (i.e., Cityscapes, KITTI 2012, and KITTI 2015), our method achieves an average BD-rate reduction of 54.61% compared to MV-HEVC and 26.81% relative to the previous state-of-the-art NSVC method LLSS [3].
Moreover, on general-scene multi-view video sequences (i.e., Kendo and Balloons), our approach attains an average BD-rate saving of 15.07% compared to MV-HEVC, and surpasses the single-view baseline DCVC-TCM* by 66.91% on average. These results highlight our model’s superior capability in leveraging disparity redundancy, particularly in indoor stereo scenes—an area where prior NSVC methods such as LSVC and LLSS have primarily demonstrated benefits only in driving scenarios.
Overall, the experimental results consistently demonstrate the superiority of our method over both traditional codecs and recent learning-based stereo video compression frameworks, achieving lower BD-rate values across all evaluated datasets.
Subjective quality comparison. Fig. 5 presents the subjective quality comparison. Compared with MV-HEVC [2], the proposed method achieves noticeably better visual quality at comparable or lower bit rates, and preserves more fine details.
IV-D Computational Complexity
@@
Table II compares the FLOPs, parameter counts (i.e., Params), and latency (i.e., Enc Speed and Dec Speed) of our model with DCVC-TCM [50], DCVC-TCM*, LSVC [4], and LLSS [3], on a $\times512$ input frame pair, following the evaluation protocol described in [3]. The FLOPs and parameter counts for LLSS are cited from its published results. We measure FLOPs, parameters, and latency for DCVC-TCM, DCVC-TCM*, LSVC, and our model using an NVIDIA RTX 4090 GPU.
As shown in Table II, DCVC-TCM*, a simplified variant of the backbone model DCVC-TCM, reduces the FLOPs from 735G to 431G, significantly lowering computational complexity. Incorporating the proposed HDC-FER and HDC-EM modules into DCVC-TCM* moderately increases the FLOPs to 536.90G.
@@
Compared to previous state-of-the-art NSVC methods such as LSVC and LLSS, our approach achieves better compression performance while maintaining significantly lower computational overhead, fewer parameters, and faster encoding and decoding speeds (compared to LSVC). These results validate the efficiency and effectiveness of the proposed modules, making our method highly suitable for deployment in resource-constrained environments.
IV-E Ablation Study
Effectiveness of HDC-FER and HDC-EM. To evaluate the individual and combined effectiveness of our proposed HDC-FER and HDC-EM modules, we conduct an ablation study based on the single-view baseline model DCVC-TCM* using the Cityscapes dataset. Specifically, the variant DCVC-TCM* + HDC-FER integrates HDC-FER modules into the feature extraction and reconstruction process of the baseline method DCVC-TCM*, enabling enhanced feature representation during the encoding and decoding stages to simultaneously compress the two views of stereo frames. DCVC-TCM* + HDC-EM replaces the original single view entropy module of DCVC-TCM* with the proposed cross-view HDC-EM entropy model, which leverages information from both views to achieve more accurate entropy estimation for each view in an autoregressive manner.
The results presented in Fig. 6 demonstrate that integrating HDC-FER alone (i.e., DCVC-TCM* + HDC-FER) achieves a BD-rate reduction of 6.03%, and substituting the entropy module with HDC-EM (i.e., DCVC-TCM* + HDC-EM) results in a BD-rate reduction of 26.59%. When both modules are simultaneously integrated into the model (i.e., DCVC-TCM* + HDC-EM + HDC-FER), the total BD-rate reduction is further improved to 32.17%, thereby confirming the significant benefits of employing cross-view information and enhanced feature representation for compression performance.
Figure 8: Ablation study on rate-distortion (RD) performance with various single-view NVC backbone networks in our framework.
Effectiveness on Implicit and Explicit Compensation in HDC. To validate the effectiveness of the implicit and explicit compensation within our HDC mechanism, we conduct an ablation study on the Cityscapes dataset, focusing on the “attention score” (i.e., implicit cross-attention) and the shift operation (i.e., explicit pixel displacement). First, removing the “attention score” from both HDC-EM and HDC-FER modules results in cross-view features being extracted solely from the shifted feature volumes, without the guidance of attention weighting. As shown in Fig. 7, this modification leads to a 13.75% increase in BD-rate, highlighting the significant role of the implicit cross-attention mechanism in improving compression efficiency. Next, to assess the contribution of explicit pixel displacement, we disable the shift operation in both modules, thereby allowing feature interaction between views only through simple channel attention. This results in a 9.08% increase in BD-rate, as illustrated in Fig. 7, demonstrating the effectiveness of explicit pixel displacement in facilitating cross-view feature extraction.
Effectiveness of Single-View NVC Backbone. To rigorously assess the effectiveness of the proposed HDC-FER and HDC-EM modules, we perform a controlled comparison on the KITTI 2012 dataset by integrating them into the same backbone as prior SOTA NSVC methods, including LSVC [4] and LLSS [3]. We also attempted to incorporate our modules with other state-of-the-art NVC backbones, such as DCVC-DC [41] and DCVC-FM [43]; however, the pretrained weights provided by these methods are highly sensitive to fine-tuning, and even using a very small learning rate results in a severe drop in performance. Therefore, for a comprehensive evaluation, we adopt the official pretrained FVC [27] model trained on the Vimeo-90K dataset [63], integrate our modules into its framework, and finetune the resulting model on Cityscapes [60], following the protocol detailed in Sec. IV-B.
As presented in Fig. 8, our method achieves superior performance over prior SOTA approaches under identical backbone settings, thereby substantiating the contribution of the proposed modules to compression efficiency. Moreover, when employing DCVC-TCM* as the backbone—which corresponds to the final configuration of our model—further performance gains are observed, reaffirming the general applicability of the proposed HDC-FER and HDC-EM modules.
TABLE III: Performance comparison of cross-view interaction strategies on BD-rate (%) and latency. The proposed method is used as the anchor for BD-rate calculation.
Interaction Strategy BD-Rate (%) Latency (ms)
Stereo Attention [13] 8.19 228.54
Mutual Attention [67] 7.03 224.45
HDC (proposed) 0.0 158.26
$\ddagger$} Latency denotes the average inference time of a forward pass, measured on a pair of $512\times 512$ stereo frames, following the setting in [3].
Cross-view Interaction Strategy Comparison. To verify the effectiveness of the proposed Hybrid Disparity Compensation (HDC) mechanism, we conduct a comprehensive comparison against state-of-the-art cross-view interaction strategies in the context of stereo image compression. Specifically, we replace the cross-view interaction component within both the HDC-EM and HDC-FER modules with the Stereo Attention mechanism [13] and the Mutual Attention mechanism [67], respectively, while keeping all other components unchanged to ensure a fair comparison. This ablation study jointly evaluates compression performance and inference latency, thereby providing a holistic assessment of the proposed HDC mechanism.
As presented in Table III, replacing the HDC mechanism with Stereo Attention results in a performance degradation of 8.19%, while substituting it with Mutual Attention leads to a 7.03% drop. Moreover, our HDC mechanism also achieves significantly lower inference latency, reducing it by approximately 30% compared to both Stereo Attention and Mutual Attention mechanisms.
V Conclusion and Future Work
In this work, we proposed a neural stereo video compression framework that incorporates a novel hybrid disparity compensation strategy to effectively reduce cross-view redundancy. By integrating explicit structural priors with implicit disparity alignment, our method achieves a synergistic balance between efficiency and robustness. Extensive experiments demonstrate that our approach achieves state-of-the-art compression performance on both automotive driving and standard indoor stereo benchmarks, outperforming existing traditional and neural stereo video compression methods. Ablation studies further validate the effectiveness and efficiency of our proposed modules. Beyond stereo video, the proposed technique shows strong potential for multi-view applications, where specifically designed cross-view interaction orders enable effective exploitation of inter-view redundancies, representing a promising direction for future research.